# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v2

@@ 8. Conclusion
In this paper, we introduce MicroScopiQ, a novel co-design technique for post-training quantization of FMs. MicroScopiQ addresses outlier-aware quantization by: a) using higher-precision MX-FP for outliers and lower-precision INTs for inliers to maintain accuracy, and b) pruning to distribute additional outlier bits, ensuring memory alignment and hardware efficiency. The MicroScopiQ accelerator, featuring multi-precision PEs and a ReCoN NoC, efficiently handles distributed outliers. MicroScopiQ is the first PTQ technique to achieve SoTA compression for both LLMs and VLMs at just $2.36$-bits, offering 3×3$\times$3 × better inference performance and 1.5×1.5$\times$1.5 × lower energy consumption than existing accelerators.
Acknowledgements
This work was supported in part by CoCoSys, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. The authors would like to thank Geonhwa Jeong for his valuable insights and feedback on this work, Zishen Wan for functional accuracy simulation, Jiawei Hu for place-and-route and Jianming Tong for his help in designing ReCoN.

### v3

@@ 8. Conclusion
In this paper, we introduce MicroScopiQ, a novel co-design technique for post-training quantization of FMs. MicroScopiQ addresses outlier-aware quantization by: a) using higher-precision MX-FP for outliers and lower-precision INTs for inliers to maintain accuracy, and b) pruning to distribute additional outlier bits, ensuring memory alignment and hardware efficiency. The MicroScopiQ accelerator, featuring multi-precision PEs and a ReCoN NoC, efficiently handles distributed outliers. MicroScopiQ is the first PTQ technique to achieve SoTA compression for both LLMs and VLMs at just $2.36$-bits, offering 3×3$\times$3 × better inference performance and 2×2$\times$2 × lower energy consumption than existing accelerators.